OpenStudy (anonymous):

I am working on this problem. I am sure it is simple, but I am just getting back into the studying thing :) The problem is: The sum of a number and its reciprocal is 10/3. What is the number? I think the formula should be x + 1/x = 10/3, but i am not sure.

OpenStudy (anonymous):

Hi! Let the number be x Reciprocal = 1/x x+1/x = 10/3 LCD=X x^2/x + 1/x = 10/3 x^2 +1 = 10x/3 x^2+1-10x/3 = 0 Now using quadratic formula find the roots! Hope it helps!

OpenStudy (anonymous):

yes that helps. thanx !

OpenStudy (anonymous):

Answer is 1/3 + 3/1 = 10/1 . Hope it helps!

OpenStudy (anonymous):

ok I am going to try to figure it out and see what happens...this is my second day back at math....so it is alot of review :)